If the line y=A+B(x-6) is perpendicular to the line y=2x and it contains the point (1,10), what are the values of A and B?
Answer by nerdybill(7384)

If the line y=A+B(x-6) is perpendicular to the line y=2x and it contains the point (1,10), what are the values of A and B?
.
slope of
y=2a
is 2
a line perpendicular must have a "negative reciprocal"
2m = -1
m = -1/2 (slope of our line)
.
y=A+B(x-6)
y=A+Bx-6B
y=Bx+A-6B
y=Bx+(A-6B)
the above is in "slope-intercept" form
where B is the slope
(A-6B) is the y-intercept
.
We found our slope above as -1/2 and we also were given one point (1,10) plug it in:
y=Bx+(A-6B)
now we know B=-1/2
10=(-1/2)(1)+(A-6(-1/2))
10=(-1/2)+(A+3)
10=(-1/2)+A+3
7=(-1/2)+A
7.5 = A
.
Answer:
B= -.5
A= 7.5